OpenStudy (anonymous):

a school has 63 students studying A,B and C. 33 study A, 25 study B, 26 study C. 10 study A and B, 9 study B and C while 8 study both A and C. Equal num study all three subjects as those who learn none of the three. How many study all 3 subjects? #help me solve this question#

OpenStudy (anonymous):

 a+b+d+e = 33 b+e+f+c = 25 d+e+f+g = 26 b+e = 10 f+e = 9 d+e = 8 e = 63 - (a+b+c+d) Solve for e..